OpenStudy (anonymous):

solve using factoring techniques x^2 + 4x + 3 = 0

OpenStudy (anonymous):

this is what I did, but I know its not right cuz the answer isn't even in the right format (x + 3) (x + 1) = 0 x +3 = 0 x+3 -3 3-0 x = -3 x + 1 = 0 x+1 -1 0 -1 x = -1 (-3, -1)

OpenStudy (anonymous):

those are correct answers

OpenStudy (anonymous):

must be they want factors, not the value of x

OpenStudy (anonymous):

some of the sample problems have an answer like "-3 or -1"... is this the same thing?

OpenStudy (anonymous):

they are the correct answers

OpenStudy (anonymous):

okay (: awesome thank you

OpenStudy (anonymous):

put those answers in the form your work shows

OpenStudy (anonymous):

sometimes as {x|x=-3, x=-1}

OpenStudy (anonymous):

I understand. I looked back up and saw "solve". When I stated it, she said it would not accept her answers, I was just throwing up a suggestion.